# Stellar properties indicating the presence of hyperons in neutron stars

### v1

@@ Abstract
We describe distinctive stellar features indicating the presence of hyperons in neutron stars. A strongly negative curvature of the mass-radius relation $R(M)$ is characteristic of hyperons, which can be determined from measurements of neutron stars with three different masses. Similarly, a reduced second derivative of the tidal deformability as function of mass $\lambda(M)$ points to hyperonic degrees of freedom in NS matter. The slopes of such curves $R(M)$ and $\lambda(M)$ can distinguish a hyperonic equation of state from purely nucleonic models if they appear increased (decreased for $\lambda(M)$) relative to the maximum mass of neutron stars.
Introduction – Non-nucleonic degrees of freedom such as hyperons are widely discussed as possible components of neutron star (NS) matter. Since the constituents of high-density matter and their interactions are not precisely known, the equation of state (EoS) of NSs is subject to uncertainties [1, 2, 3, 4, 5, 6, 7]. The EoS of cold NSs in beta equilibrium is a unique relation between pressure and energy density that unambiguously determines the stellar structure like the mass-radius relation or the tidal deformability of stars in binaries through the relativistic stellar structure equations [8, 9, 10, 11, 12]. By measuring stellar parameters of NSs it is thus possible to probe the EoS and learn about the fundamental building blocks of high-density matter, e.g. the presence of hyperons.
@@
The capabilities of NS observations to determine stellar properties are continuously growing for instance with the X-ray timing measurements of NICER [34, 35, 36, 37] or with gravitational-wave observatories [38, 39, 40, 41], and several new instruments are projected to advance these efforts in the next decade [42, 43, 44, 45]. But the current lack of a distinct feature of hyperons imprinted on the stellar structure parameters implies that even precise measurements of NS properties would not allow to infer the hyperon content of NSs and solve the “hyperon puzzle” (cf. [14, 15, 16, 17, 18, 19, 6, 20]).
@@
In this paper we propose to consider the curvature and the second derivative with respect to mass of the mass-radius curve and the mass-tidal deformability relation to identify the presence of hyperons in NSs. Those parameters quantify the bending towards smaller NS radii with increasing mass and are found to be characteristically lowered if hyperons are present. We also find that the slope of the aforementioned relations at a fixed mass compared to the maximum mass of NSs can indicate hyperonic degrees of freedom. Observationally, these characteristics are accessible by considering several measurements of NSs with different masses.
@@
EoS sample – We base our study on a very large set of microphysical EoS models for matter in beta equilibrium and at zero temperature. This sample includes 46 hyperonic models and 247 purely nucleonic EoSs (see Supplemental Material). These tables stem from three sources, to which we refer for more details. One subset consists of various models which we previously used in NS merger simulations [46, 47] and which were obtained through private communication or various webpages [48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83]. Another larger set of models is taken from the Compose repository [13], where we include all baryonic NS matter models with a maximum mass of at least 1.9 $M_{\odot}$ [84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129, 63, 130, 131, 132, 133, 134, 135, 136, 137, 138, 139]. We supplement our sample with a selection of models for purely nucleonic [140, 141] and hyperonic [142] matter constructed from covariant density functional theory. Overall, our sample represents a large variety of models which are based on different theoretical frameworks (e.g. non-relativistic and relativistic density functional methods) and adjusted to different experimental data (properties of nuclear matter and nuclei). We include parametric models only if they are fitted to reproduce actual microphysical models and we refrain from using agnostic approaches to the EoS since they do not inform about the composition and thus the potential presence of hyperons. All EoS models reach a maximum NS mass of at least 1.9 $M_{\odot}$, which is somewhat below the limit imposed by heavy pulsars [143, 144, 145, 146, 147], but it is advantageous to slightly enlarge the range of possible models (only 17 models have $M_{\mathrm{max}}<2.0~{}M_{\odot}$). Otherwise we do not employ any additional selection criteria for the EoS sample albeit various constraints might have already been imposed during the development of certain EoS models.
@@
Radius curves – For all EoSs we compute the circumferential radii $R$ and tidal deformability $\lambda$ as function of gravitational mass $M$ using the extended set of relativistic stellar structure equations [8, 9, 10, 11, 12]. $\lambda$ is given by $\frac{2}{3}k_{2}R^{5}$ with the tidal Love number $k_{2}$. All curves are provided in the Supplemental Material. In mass-radius relations (Fig. 1) hyperonic models often exhibit the tendency to more strongly bend over to smaller radii with increasing mass. We intend to assess and quantify this effect as a characteristic indicator of hyperons in NSs.
@@
The small curvature for hyperonic EoSs is a consequence of the softening of the EoS at densities where hyperons occur. The mass where $\kappa_{R}$ starts to significantly drop approximately corresponds to the onset density of hyperon production. A strongly negative curvature is found for most hyperonic EoSs of our sample although there are a few exceptions (7 out of 49 hyperonic models do not reach below $\kappa_{R}=-1.5$).
@@
This finding suggests to measure $\kappa_{R}$ in the range $1~{}M_{\odot}\lesssim M\lesssim 2~{}M_{\odot}$ to discriminate a hyperonic EoS from purely nucleonic EoSs. In particular in the range $1.4~{}M_{\odot}\lesssim M\lesssim 2.0~{}M_{\odot}$ there exist no nucleonic model reaching below $\kappa_{R}=-2.5$. The smallest $\kappa_{R}$ reached by nucleonic models depends on mass, and in fact the bulk of the nucleonic models does not reach below $\kappa_{R}=-1.5$ for $M<2.0~{}M_{\odot}$ (below $\kappa_{R}=-1.0$ for $M<1.6~{}M_{\odot}$). The distinction between hyperonic models and nucleonic models could even become more pronounced by checking those nucleonic models with a particularly small $\kappa_{R}$ and whether those can be disregarded based on physics arguments. Note that $\kappa_{R}$ not dropping below -1.5 does however not exclude the presence of hyperons.
@@
The accuracy of about 0.1 km is not yet available in current measurements, although on longer terms an order of magnitude improvement of the error might not be completely unrealistic [148, 149, 150, 151, 152, 153]. We also note that statistical approaches and the inclusion of more than three measurements may improve the determination of $\kappa_{R}$. The exact requirements for an identification of hyperons through $\kappa_{R}$ also depend on the considered reference mass. For instance, at masses above 1.6 $M_{\odot}$ the differences in $\kappa_{R}$ between hyperonic and nucleonic models become more pronounced (Fig. 2) but obtaining measurements with sufficiently large $\Delta M$ might be less likely as one approaches $M_{\mathrm{max}}$.
$\frac{d^{2}R}{dM^{2}}$ at a given mass $1.5~{}M_{\odot}\lesssim M_{\mathrm{ref}}\lesssim 2.0~{}M_{\odot}$ versus $R(M_{\mathrm{ref}})$ may also be an interesting indicator for hyperons, where several hyperonic models tend to lie at lower values than purely nucleonic EoSs (similar to Fig. 3 but less pronounced).
Tidal deformability curves – We now discuss the tidal deformability as indicator of hyperonic EoSs. We remark that the occurrence of features of hyperonic models in the first and second derivative with respect to mass and the curvature $\kappa$ depends on whether one considers $\lambda(M)$ or $\Lambda(M)=\lambda(M)/M^{5}$ because the specific effects of hyperons can cancel each other in different terms.
@@
It is similarly possible to consider $\Lambda(M)$, where hyperonic models do not stand out prominently in relations like $\frac{d^{2}\Lambda}{dM^{2}}(M)$ or $\kappa_{\Lambda}(M)$. Evaluating those quantities at fixed reference mass $1.6~{}M_{\odot}\lesssim M_{\mathrm{ref}}\lesssim 2.0~{}M_{\odot}$ versus $\Lambda(M_{\mathrm{ref}})$, some hyperonic models yield a characteristically reduced $\kappa_{\Lambda}(M_{\mathrm{ref}})$ compared to purely nucleonic EoSs, whereas $\frac{d^{2}\Lambda}{dM^{2}}(M_{\mathrm{ref}})$ is elevated (but some nucleonic models are similarly enhanced). For those hyperonic models, measurements of $\Lambda$ within a few per cent are sufficient to identify this signature of hyperons. However, in comparison to $\frac{d^{2}\lambda}{dM^{2}}$ the signatures are not as pronounced and especially for $\frac{d^{2}\Lambda}{dM^{2}}$ several purely nucleonic models overlap.
Figure 6: Slope of $R(M)$ (top) and $\Lambda(M)$ (bottom) at a fixed mass of $M_{\mathrm{ref}}=1.6~{}M_{\odot}$ as function of the maximum mass of NS. Symbols and colors have the same meaning as in Fig. 3.
Slopes – Finally, we highlight the slopes $\frac{dR}{dM}$ and $\frac{d\Lambda}{dM}$ as indicator for hyperons [154, 155, 156, 157, 158, 159, 21]. Reference [21] argued that $\frac{dR}{dM}<0$ in the range $1.2~{}M_{\odot}\leq M\leq 1.4~{}M_{\odot}$ would favor a nucleonic model employing statistical considerations without noting a distinctive feature for hyperonic EoSs. The slope of $R(M)$ or similarly of $\lambda(M)$ and $\Lambda(M)$ at a fixed mass on its own is not very conclusive with regards to the presence of hyperons because nucleonic models overlap over the whole range of hyperonic models (see Supplemental Material; with $\frac{d\lambda}{dM}$ appearing slightly more promising).
We however point out that comparing the slopes at a fixed mass to the maximum mass, specific hyperonic models stand out clearly. Figure 6 shows $\frac{dR}{dM}$ and $\frac{d\Lambda}{dM}$ evaluated at $M_{\mathrm{ref}}=1.6~{}M_{\odot}$ as function of $M_{\mathrm{max}}$ demonstrating that an increased (decreased) slope in $\frac{dR}{dM}$ ($\frac{d\Lambda}{dM}$) compared to $M_{\mathrm{max}}$ is a distinctive feature of hyperonic models. For $\frac{d\Lambda}{dM}$ there are however two nucleonic models with similarly small slope. For $\frac{d\lambda}{dM}$ hyperonic EoSs do not occur distinguished. Similar findings hold for $1.2~{}M_{\odot}\lesssim M_{\mathrm{ref}}\lesssim 1.8~{}M_{\odot}$. At first glance, an increased slope of $R(M)$ may seem counterintuitive as indicator of hyperons as one would associate it with a certain stiffness of the EoS. In those models hyperons soften the EoS at higher densities and thus reduce $M_{\mathrm{max}}$ more strongly than any purely nucleonic model with the same slope.
@@
Discussion and outlook – We describe stellar features which allow to quantitatively distinguish a hyperonic EoS from purely nucleonic EoSs. Requirements for corresponding measurements are generally challenging but not completely out of reach [148, 149, 150, 151, 152, 153]. In particular, the required precision for measurements of the tidal deformability of a few per cent might be feasible with the next generation of GW detectors [42, 44] although this precision has to be achieved for high-mass NSs. Future work should thus explore the capabilities of new instruments [42, 43, 44, 45] and the potential of statistical methods including for instance more than three measurements.
@@ Acknowledgements.
We thank David Blaschke, Gabriel Martinez-Pinedo and Angels Ramos for helpful discussions. A.B. and A.N. acknowledge support by the European Union through ERC Synergy Grant HeavyMetal no. 101071865. A.B. acknowledges support by Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through Project-ID 279384907 – SFB 1245 (subproject B07). G.L acknowledges support by the Klaus Tschira Foundation. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 101034371. PC acknowledges the support from the European Union’s HORIZON MSCA-2022-PF-01-01 Programme under Grant Agreement No. 101109652, project ProMatEx-NS. C.M. acknowledges partial support from the Fonds de la Recherche Scientifique (FNRS, Belgium) and the Research Foundation Flanders (FWO, Belgium) under the EOS Project nr O022818F and O000422. M.O. acknowledges financial support of the Agence Nationale de la Recherche (ANR) under contract ANR-22-CE31-0001-01. L.T. acknowledges support from the program Unidad de Excelencia María de Maeztu CEX2020-001058-M, from the project PID2022-139427NB-I00 financed by the Spanish MCIN/AEI/10.13039/501100011033/FEDER, UE (FSE+), from the Grant CIPROM 2023/59 of Generalitat Valenciana and from CRC-TR 211 ’Strong-interaction matter under extreme conditions’, Project No. 315477589 - TRR 211 by the Deutsche Forschungsgemeinshaft. S.G. and N.C acknowledge financial support from F.R.S.-FNRS (Belgium). This work was supported by the Fonds de la Recherche Scientifique - FNRS and the Fonds Wetenschappelijk Onderzoek - Vlaanderen (FWO) under the EOS Project No O000422.
I Supplemental Material

### v2

@@ Abstract
We describe distinctive stellar features indicating the presence of hyperons in neutron stars as compared to purely nucleonic systems. A strongly negative curvature of the mass-radius relation $R(M)$ is characteristic of hyperons, which can be determined from measurements of neutron stars with three different masses. Similarly, a reduced second derivative of the tidal deformability as function of mass $\lambda(M)$ points to hyperonic degrees of freedom in NS matter. The slopes of such curves $R(M)$ and $\lambda(M)$ can distinguish a hyperonic equation of state from purely nucleonic models if they appear increased (decreased for $\lambda(M)$) relative to the maximum mass of neutron stars.
I Introduction
Non-nucleonic degrees of freedom such as hyperons are widely discussed as possible components of neutron star (NS) matter. Since the constituents of high-density matter and their interactions are not precisely known, the equation of state (EoS) of NSs is subject to uncertainties [1, 2, 3, 4, 5, 6, 7]. The EoS of cold NSs in beta equilibrium is a unique relation between pressure and energy density that unambiguously determines the stellar structure like the mass-radius relation or the tidal deformability of stars in binaries through the relativistic stellar structure equations [8, 9, 10, 11, 12]. By measuring stellar parameters of NSs it is thus possible to probe the EoS and learn about the fundamental building blocks of high-density matter, e.g. the presence of hyperons. These efforts are linked to numerous theoretical and experimental studies to understand the interactions between hyperons and nucleons [13, 14, 15].
@@ I Introduction
The capabilities of NS observations to determine stellar properties are continuously growing for instance with the X-ray timing measurements of NICER [37, 38, 39, 40] or with gravitational-wave observatories [41, 42, 43, 44], and several new instruments are projected to advance these efforts in the next decade [45, 46, 47, 48]. But the current lack of a distinct feature of hyperons imprinted on the stellar structure parameters implies that even precise measurements of NS properties would not allow to infer the hyperon content of NSs and solve the “hyperon puzzle” (cf. [49, 17, 18, 19, 20, 21, 22, 6, 23, 14, 50, 15]).
@@
In this paper we propose to consider the curvature and the second derivative with respect to mass of the mass-radius curve and the mass-tidal deformability relation to identify the presence of hyperons in NSs or, generally, non-nucleonic degrees of freedom. Those parameters quantify the bending towards smaller NS radii with increasing mass and are found to be characteristically lowered if hyperons are present. We also find that the slope of the aforementioned relations at a fixed mass compared to the maximum mass of NSs can indicate hyperonic degrees of freedom. Observationally, these characteristics are accessible by considering several measurements of NSs with different masses. The stellar structure does not give a handle on which mechanisms and degrees of freedom are responsible for softening the EoS. Thus, ultimately, the properties discussed here can only indicate that non-nucleonic degrees of freedom soften the EoS so strongly that it is incompatible with purely nucleonic models. Non-nucleonic degrees of freedom appearing via a first-order phase transition (such as quarks in many models) result in a characteristic kink in the mass-radius relation, which has been discussed as potential signature for many years [51, 1, 2, 6], and thus identifying a quantitative signature for a smoother occurrence of non-nucleonic degrees of freedom remains a major challenge.
II EoS sample
@@
We base our study on a very large set of microphysical EoS models for matter in beta equilibrium and at zero temperature. This sample includes 47 hyperonic models and 248 purely nucleonic EoSs (see appendix). These tables stem from three sources, to which we refer for more details. One subset consists of various models which we previously used in NS merger simulations [52, 53] and which were obtained through private communication or various webpages [54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91]. Another larger set of models is taken from the Compose repository [16], where we include all baryonic NS matter models with a maximum mass of at least 1.9 $M_{\odot}$ [92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136, 137, 70, 138, 139, 140, 141, 142, 143, 144, 145, 146, 147, 148, 149, 150, 151, 152, 153, 154]. We supplement our sample with a selection of models for purely nucleonic [155, 156] and hyperonic [157, 158] matter constructed from covariant density functional theory. Overall, our sample represents a large variety of models which are based on different theoretical frameworks (e.g. non-relativistic and relativistic density functional methods) and adjusted to different experimental data (properties of nuclear matter and nuclei). We include parametric models only if they are fitted to reproduce actual microphysical models and we refrain from using agnostic approaches to the EoS since they do not inform about the composition and thus the potential presence of hyperons. All EoS models reach a maximum NS mass of at least 1.9 $M_{\odot}$, which is somewhat below the limit imposed by heavy pulsars [159, 160, 161, 162, 163], but it is advantageous to slightly enlarge the range of possible models (only 9 hyperonic and 8 purely nucleonic models have $M_{\mathrm{max}}<2.0~M_{\odot}$). Otherwise we do not employ any additional selection criteria for the EoS sample albeit various constraints might have already been imposed during the development of certain EoS models.
@@
sectionRadius curves For all EoSs we compute the circumferential radii $R$ and tidal deformability $\lambda$ as function of gravitational mass $M$ using the extended set of relativistic stellar structure equations [8, 9, 10, 11, 12]. $\lambda$ is given by $\frac{2}{3}k_{2}R^{5}$ with the tidal Love number $k_{2}$. All curves are provided in the appendix. In mass-radius relations (Fig. 1) hyperonic models often exhibit the tendency to more strongly bend over to smaller radii with increasing mass. We intend to assess and quantify this effect as a characteristic indicator of hyperons in NSs.
@@
The small curvature for hyperonic EoSs is a consequence of the softening of the EoS at densities where hyperons occur. The mass where $\kappa_{R}$ starts to significantly drop approximately corresponds to the onset density of hyperon production. A strongly negative curvature is found for most hyperonic EoSs of our sample although there are a few exceptions (7 out of 47 hyperonic models do not reach below $\kappa_{R}=-1.5$). Considering the ten hyperonic models with the smallest minima of the curvature, two models have $M_{\mathrm{max}}<2.0\,M_{\odot}$ while eight EoSs are compatible with the constraints from pulsar measurements.
@@
This finding suggests to measure $\kappa_{R}$ in the range $1~M_{\odot}\lesssim M\lesssim 2~M_{\odot}$ to discriminate a hyperonic EoS from purely nucleonic EoSs. In particular in the range $1.4~M_{\odot}\lesssim M\lesssim 2.0~M_{\odot}$ there exist no nucleonic model reaching below $\kappa_{R}=-2.5$. The smallest $\kappa_{R}$ reached by nucleonic models depends on mass, and in fact the bulk of the nucleonic models does not reach below $\kappa_{R}=-1.5$ for $M<2.0~M_{\odot}$ (below $\kappa_{R}=-1.0$ for $M<1.6~M_{\odot}$). Note that $\kappa_{R}$ not dropping below -1.5 does however not exclude the presence of hyperons.
@@
The accuracy of about 0.1 km is not yet available in current measurements, although on longer terms an order of magnitude improvement of the error might not be completely unrealistic [164, 165, 166, 167, 168, 169, 170]. We also note that statistical approaches and the inclusion of more than three measurements may improve the determination of $\kappa_{R}$. The exact requirements for an identification of hyperons through $\kappa_{R}$ also depend on the considered reference mass. For instance, at masses above 1.6 $M_{\odot}$ the differences in $\kappa_{R}$ between hyperonic and nucleonic models become more pronounced (Fig. 2) but obtaining measurements with sufficiently large $\Delta M$ might be less likely as one approaches $M_{\mathrm{max}}$.
$\frac{d^{2}R}{dM^{2}}$ at a given mass $1.5~M_{\odot}\lesssim M_{\mathrm{ref}}\lesssim 2.0~M_{\odot}$ versus $R(M_{\mathrm{ref}})$ may also be an interesting indicator for hyperons, where several hyperonic models tend to lie at lower values than purely nucleonic EoSs (similar to Fig. 3 but less pronounced, see appendix).
III Tidal deformability curves
We now discuss the tidal deformability as indicator of hyperonic EoSs. We remark that the occurrence of features of hyperonic models in the first and second derivative with respect to mass and the curvature $\kappa$ depends on whether one considers $\lambda(M)$ or $\Lambda(M)=\lambda(M)/M^{5}$ because the specific effects of hyperons can cancel each other in different terms.
@@ III Tidal deformability curves
It is similarly possible to consider $\Lambda(M)$, where hyperonic models do not stand out prominently in relations like $\frac{d^{2}\Lambda}{dM^{2}}(M)$ or $\kappa_{\Lambda}(M)$. Evaluating those quantities at fixed reference mass $1.6~M_{\odot}\lesssim M_{\mathrm{ref}}\lesssim 2.0~M_{\odot}$ versus $\Lambda(M_{\mathrm{ref}})$, some hyperonic models yield a characteristically reduced $\kappa_{\Lambda}(M_{\mathrm{ref}})$ compared to purely nucleonic EoSs, whereas $\frac{d^{2}\Lambda}{dM^{2}}(M_{\mathrm{ref}})$ is elevated (but some nucleonic models are similarly enhanced). For those hyperonic models, measurements of $\Lambda$ within a few per cent are sufficient to identify this signature of hyperons. However, in comparison to $\frac{d^{2}\lambda}{dM^{2}}$ the signatures are not as pronounced and especially for $\frac{d^{2}\Lambda}{dM^{2}}$ several purely nucleonic models overlap (see appendix).
Figure 6: Slope of $R(M)$ (top) and $\Lambda(M)$ (bottom) at a fixed mass of $M_{\mathrm{ref}}=1.6~M_{\odot}$ as function of the maximum mass of NS. Symbols and colors have the same meaning as in Fig. 3 with error bars computed for the values mentioned in the main text
IV Slopes
Finally, we highlight the slopes $\frac{dR}{dM}$ and $\frac{d\Lambda}{dM}$ as indicator for hyperons (cf. [171, 172, 173, 174, 175, 176, 24, 177, 178]. Reference [24] argued that $\frac{dR}{dM}<0$ in the range $1.2~M_{\odot}\leq M\leq 1.4~M_{\odot}$ would favor a nucleonic model employing statistical considerations without noting a distinctive feature for hyperonic EoSs. The slope of $R(M)$ or similarly of $\lambda(M)$ and $\Lambda(M)$ at a fixed mass on its own is not very conclusive with regards to the presence of hyperons because nucleonic models overlap over the whole range of hyperonic models (see appendix; with $\frac{d\lambda}{dM}$ appearing slightly more promising).
We however point out that comparing the slopes at a fixed mass to the maximum mass, some specific hyperonic models stand out clearly. Figure 6 shows $\frac{dR}{dM}$ and $\frac{d\Lambda}{dM}$ evaluated at $M_{\mathrm{ref}}=1.6~M_{\odot}$ as function of $M_{\mathrm{max}}$ demonstrating that an increased (decreased) slope in $\frac{dR}{dM}$ ($\frac{d\Lambda}{dM}$) compared to $M_{\mathrm{max}}$ is a distinctive feature of hyperonic models. For $\frac{d\Lambda}{dM}$ there are however two nucleonic models with similarly small slope. For $\frac{d\lambda}{dM}$ hyperonic EoSs do not occur distinguished. Similar findings hold for $1.2~M_{\odot}\lesssim M_{\mathrm{ref}}\lesssim 1.8~M_{\odot}$. At first glance, an increased slope of $R(M)$ may seem counterintuitive as indicator of hyperons as one would associate it with a certain stiffness of the EoS. In those models hyperons soften the EoS at higher densities and thus reduce $M_{\mathrm{max}}$ more strongly than any purely nucleonic model with the same slope.
@@ IV Slopes
V Discussion and outlook
We describe stellar features which allow to quantitatively distinguish a hyperonic EoS from purely nucleonic EoSs. Requirements for corresponding measurements are generally challenging but not completely out of reach [164, 165, 166, 167, 168, 169, 170]. In particular, the required precision for measurements of the tidal deformability of a few per cent might be feasible with the next generation of GW detectors [45, 47] although this precision has to be achieved for high-mass NSs. Future work should thus explore the capabilities of new instruments [45, 46, 47, 48] and the potential of statistical methods including for instance more than three measurements.
@@ Acknowledgements.
We thank David Blaschke, Gabriel Martinez-Pinedo and Angels Ramos for helpful discussions, and Hui Tong for providing EoS tables. A.B. and A.N. acknowledge support by the European Union through ERC Synergy Grant HeavyMetal no. 101071865. A.B. acknowledges support by Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through Project-ID 279384907 – SFB 1245 (subproject B07). G.L acknowledges support by the Klaus Tschira Foundation. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 101034371. PC acknowledges the support from the European Union’s HORIZON MSCA-2022-PF-01-01 Programme under Grant Agreement No. 101109652, project ProMatEx-NS. C.M. acknowledges partial support from the Fonds de la Recherche Scientifique (FNRS, Belgium) and the Research Foundation Flanders (FWO, Belgium) under the EOS Project nr O022818F and O000422. M.O. acknowledges financial support of the Agence Nationale de la Recherche (ANR) under contract ANR-22-CE31-0001-01. L.T. acknowledges support from the program Unidad de Excelencia María de Maeztu CEX2020-001058-M, from the project PID2022-139427NB-I00 financed by the Spanish MCIN/AEI/10.13039/501100011033/FEDER, UE (FSE+), from the Grant CIPROM 2023/59 of Generalitat Valenciana and from CRC-TR 211 ’Strong-interaction matter under extreme conditions’, Project No. 315477589 - TRR 211 by the Deutsche Forschungsgemeinshaft. S.G. and N.C acknowledge financial support from F.R.S.-FNRS (Belgium). This work was supported by the Fonds de la Recherche Scientifique - FNRS and the Fonds Wetenschappelijk Onderzoek - Vlaanderen (FWO) under the EOS Project No O000422.
Appendix A EoS models and additional plots